OpenStudy (anonymous):

The sum of the squares of two consecutive positive numbers is 85. What is the smaller number?

OpenStudy (anonymous):

try \(3^2+4^2\) and if that is too small, try \(4^2+5^2\) or \(5^2+6^2\)

OpenStudy (anonymous):

@ParthKohli sum of squares

OpenStudy (anonymous):

Let one number be: x Next number will be: (x + 1) So, \[x^2 + (x +1)^2 = 85\] \[(a + b)^2 = a^2 + b^2 + 2ab\] Use this formula to solve this above equation...

OpenStudy (anonymous):

@sabimaru sum of squares

Parth (parthkohli):

Thanks.. will get my eyes checked

Parth (parthkohli):

Thing is that the next time I'd go there in October

OpenStudy (anonymous):

\[6^2+7^2=85\] why will no one guess and check? it take like 5 seconds with a calculator

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6^2%2B7^2

Parth (parthkohli):

@calculator73 okay